OpenStudy (anonymous):

what is the integration of( x+2)/(x+1)dx ?

OpenStudy (cwrw238):

divide firsr

OpenStudy (anonymous):

We can either divide it out or we can do something along the lines of this: let u = x+1 du = dx. Becauseu = x+1, the numerator is just u + 1 \[\int\limits_{}^{}\frac{ u + 1 }{ u }du \implies \int\limits_{}^{}(\frac{ u }{ u }+ \frac{ 1 }{ u })du \implies \int\limits_{}^{}(1 + \frac{ 1 }{ u })du\]